OpenStudy (hellokitty17):

A right triangle has base (x - 6) units and height (6x - 12) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points)

OpenStudy (hoothoot):

correlation coefficient is 1 slope is 5/10 which reduces to 1/2 not sure about the 3rd one

OpenStudy (jack1):



oregonduck (oregonduck):

@HootHoot help plz

OpenStudy (phi):

do you know the Pythagorean Theorem? (it's both famous and useful for this problem)

OpenStudy (hoothoot):

OK ILL GET MY BRO ON IT

oregonduck (oregonduck):

kk

OpenStudy (hoothoot):

ok i sent the question to em he is working on it

OpenStudy (hoothoot):

I think it has to do with A^2 + B^2 = C^2 where C^2 is your hypotenuse. (x-6)^2 + (6x - 12)^2 = c^2.ok he belives it like this

OpenStudy (hoothoot):

whats epic

OpenStudy (anonymous):

i tinhk hoot hoot is correct

OpenStudy (anonymous):

do your foiling and stuff

OpenStudy (anonymous):

f= first, o=outside, i=inside, l=lasts

OpenStudy (hellokitty17):

we know that

OpenStudy (hoothoot):

it shouldnt be that much harder from here

OpenStudy (anonymous):

(6x-13)^2 = (6x-12)(6x-12)

OpenStudy (anonymous):

do the same thing for the base

oregonduck (oregonduck):

write it out like Part a... Part b... and part c...

OpenStudy (hellokitty17):

@jcoury

oregonduck (oregonduck):

@jcoury

oregonduck (oregonduck):

Part A (6x-12)^2 + (x - 6)^2 (36x^2 - 144x +144) + (x^2 - 12x + 36)

oregonduck (oregonduck):

\[37^{2}-156x+180\]

oregonduck (oregonduck):

Part B: The area A of a triangle is A=(1/2)bh where b is base and h is height. Refer to diagram above A=(1/2)(x-6)(6x-12) simplify by taking 1/2 of the height h getting (3x-6) now multiply (x-6)(3x-6) getting: 3x^2 - 24x + 36 sq units

oregonduck (oregonduck):

Part C. Just show that the multiplication of two polynomials (x-6)(3x-6) results in a polynomial 3x^2 - 24x + 36 as a result of the distributive process and combing of similar terms.

OpenStudy (anonymous):

nice XD

oregonduck (oregonduck):

XD lol

OpenStudy (anonymous):

@OregonDuck can u help me real quick

oregonduck (oregonduck):

help with part c. sure

oregonduck (oregonduck):

tag me